A mixture of He, N2, and Ar has a pressure of 25.6 atm at 28.0 °C. If the partial pressure of He is 965 torr and that of Ar is 2969 mm Hg, what is the partial pressure of N2?
Answer #1

Partial pressure of He = 965/760 => 1.27 atm

Partial pressure of Ar = 2969/760 => 3.91 atm

P​​​​​​N2​​​​ = 25.6-(1.27+3.91)

=> 20.4 atm